# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v2

@@ IV Discussion
The booming field of machine learning and artificial intelligence is transforming the field of biomolecular modeling at a swift pace. Recent advancements in generative models, such as AlphaFoldJumper et al. (2021), RosettaFoldBaek et al. (2023) and others, have heralded a new era in the accurate prediction of structural ensembles. Leveraging the predictive power of these models as structural priors is expected to help refine ensemble predictions when integrated with similar algorithms to BICePsBrotzakis, Zhang, and Vendruscolo (2023). Here, we have demonstrated that structural ensembles generated from RosettaFold2 (RF2) can be reweighted to better align with experimental measurements, while simultaneously refining Karplus parameters. Validation of these parameters by the BICePs score and other statistics demonstrates improved accuracy across a varity of structural ensembles of ubiquitin.
@@
This representation ensures continuity across the periodic boundary at ±$\pm\pi$ and provides a smooth input feature space for learning.
@@
The network was trained using ADAM optimization for 2000 epochs with a learning rate of $10^{-3}$, using the Good-Bad likelihood model with the BICePs score as the loss function. The toy dataset consisted of 5 conformational states, each with 60 synthetic $J$-coupling observables using the Karplus relation $J(\phi)=A\cos^{2}(\phi)+B\cos(\phi)+C$, with the ground truth coefficients $A=6.51$, $B=-1.76$, and $C=1.60$. Our training dataset consisted of 300 $\phi$-angles as NN inputs.
First, we test the performance without adding random or systematic noise to the synthetic experimental J-coupling observables. Figure S16a shows the convergence behavior of the BICePs score across five independent training runs, where a plateau is reached around step 1500. Figure S16b shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). For comparison, we also include the result of singular value decomposition (SVD) applied directly to the Karplus form under the assumption that backbone $\phi$ angles are known precisely, as would be the case with a high-resolution crystal structure. The parameters from SVD were determined to be $A=6.47\pm 0.002$, $B=-1.75\pm 0.001$, and $C=1.62\pm 0.001$.
Next, we evaluate the performance when adding random or systematic noise to the synthetic experimental $J$-coupling observables ($\sigma_{\text{data}}=0.79$ Hz). Figure S17a–b shows the convergence behavior of the BICePs score and gradient norm across five independent training runs, where convergence begins around 1000 epochs. The BICePs score was determined to be $f_{\xi=0\rightarrow 1}=62.29$ nats. Figure S17c shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). Comparison against SVD illustrates that BICePs, when used to train neural networks from conformational ensembles alone, recovers predictive performance comparable to analytical fitting based on ground-truth dihedral angles, even in the presence of random and systematic errors.
@@
Models have limited predictive reliability in regions lacking sufficient training data. A central advantage of our framework lies in its ability to incorporate physically relevant information through the choice of input features or explicit regularization. In our example using the toy model system, we employed the backbone dihedral angle $\phi$ as the sole input feature (projected to the unit circle to maintain periodicity). However, neural network inputs can encompass a broader range of features for complex observables such as chemical shifts, including quantum calculations, partial charges, descriptors of local atomic environments, interatomic distances, hydrogen bonding patterns, solvent accessibility, or torsional strain energies. These diverse features encode crucial structural and electronic context and can be precomputed from molecular simulations or quantum calculations before training.
@@
We did not impose physical constraints or priors when training our model to predict $J$-coupling constants. Users might explicitly incorporate such physical priors or constraints through architectural decisions, regularization terms, or physically-informed input features, especially critical when dealing with sparse or unevenly distributed training data. For instance, in our experiments, neural networks employing Rectified Linear Unit (ReLU) activations struggled to accurately predict $J$-couplings at backbone dihedral angles greater than $+90^{\circ}$. In contrast, GELU activations provided smoother and more reliable extrapolation in these underrepresented regions. Thus, careful selection of activation functions and comprehensive input feature representations significantly enhances predictive accuracy across the full conformational landscape.
@@ V Conclusion
The BICePs algorithm is freely available at github.com/vvoelz/biceps. All scripts, Jupyter notebook examples and analysis can be found here. Detailed documentation and tutorials can be found here: biceps.readthedocs.io. For any issues or questions, please submit the request on GitHub.
@@ Acknowledgements
The authors acknowledge that substantial content in this manuscript has also been included in the doctoral dissertation of RMR.Raddi (2024) RMR, TM and VAV are supported by National Institutes of Health grant R01GM123296. This research includes calculations carried out on HPC resources supported in part by the National Science Foundation through major research instrumentation grant number 1625061 and by the US Army Research Laboratory under contract number W911NF-16-2-0189.
@@
$\frac{\partial^{2}u}{\partial\theta_{m}^{2}}=N\left[\sum_{j=1}^{N_{d}}\left(-\frac{\partial^{2}g_{j}(X,\theta)}{\partial\theta_{m}^{2}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right.\right.$ (S2)
@@ Methods
Figure S2: BICePs predicted forward model parameters in the presence of random and systematic error ($\sigma_{data}=0.471$ Hz) for a toy model system. (a) Karplus curves predicted for SVD (orange) and BICePs (blue), where the "true" (black dashed line) parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$. The extracted parameters from the SVD fitting were found to be $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$ and the BICePs was $\{A=6.6\pm 0.037,B=-1.8\pm 0.016,C=1.5\pm 0.027\}$, averaged over three independent chains. The uncertainty is represented by the thickness of the curves. For the BICePs calculation, we used the Good-Bad likelihood model with 32 replicas and burned for 20k steps, followed by 50k steps of additional sampling. Red dots correspond to the syntetic experimental J-coupling data points. (b) The marginal posterior distribution of uncertainty $p(\sigma_{J})$. The maximum a posteriori was determined to be $\hat{\sigma}_{J}=0.272$ Hz, and the a posteriori variance scaling parameter $\hat{\varphi}_{J}=1.98$. (c) Landscapes of the BICePs score, $f$ for pairs of Karplus coefficients. (d) The marginal posterior distribution of Karplus coefficients.
Figure S3: The Student’s model gives similar performance to the Good-Bad model. Karplus coefficients predicted using the Student’s likeliood model ($\{A=6.8\pm 0.033,B=-1.9\pm 0.025,C=1.4\pm 0.033\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$ Hz). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2.
@@
Figure S10: Traces of sampled Karplus coefficients for the 1d3z ensemble over 50k steps of MCMC, post-burn. BICePs calculations used the Good-Bad model with 32 replicas. Traces display low variance with no jumps, which demonstrates converged samples
@@
Figure S16: BICePs trains NN parameters for $J$-coupling predictions using a toy model system. (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1500 training steps. (b) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters ${A=6.51,B=-1.76,C=1.60}$. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters ${A=6.47\pm 0.002,B=-1.75\pm 0.001,C=1.62\pm 0.001}$. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. No random or systematic noise was added to the synthetic data. Red dots correspond to the syntetic experimental J-coupling data points.
Figure S17: BICePs trains NN parameters for $J$-coupling predictions using a toy model system in the presence of random and systematic error ($\sigma_{\text{data}}=0.79$ Hz). (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1000 training steps. (b) The average gradient norm across the five chains during training. (c) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters ${A=6.51,B=-1.76,C=1.60}$. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters ${A=6.39\pm 0.13,B=-1.55\pm 0.06,C=1.94\pm 0.10}$. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. Red dots correspond to the syntetic experimental $J$-coupling data points.

### v3

@@ IV Discussion
The booming field of machine learning and artificial intelligence is transforming the field of biomolecular modeling at a swift pace. Recent advancements in generative models, such as AlphaFoldJumper et al. (2021), RosettaFoldBaek et al. (2023) and others, have heralded a new era in the accurate prediction of structural ensembles. Leveraging the predictive power of these models as structural priors is expected to help refine ensemble predictions when integrated with similar algorithms to BICePsBrotzakis et al. (2025). Here, we have demonstrated that structural ensembles generated from RosettaFold2 (RF2) can be reweighted to better align with experimental measurements, while simultaneously refining Karplus parameters. Validation of these parameters by the BICePs score and other statistics demonstrates improved accuracy across a variety of structural ensembles of ubiquitin.
@@
Comparison of joint posterior sampling and variational minimization for FM parameterization.
Two strategies can be used for forward model parameter optimization in BICePs: joint posterior sampling and variational minimization. Joint posterior sampling, which we recommend, is efficient with modern gradient-based samplers such as Hamiltonian Monte Carlo (HMC), provides direct estimates of parameter uncertainties and covariances, and is fast in practice (50,000 steps in 15 seconds on a MacBook M1 Pro). Variational minimization of the BICePs score can be advantageous when handling many parameters and is straightforward to adapt for different forward models, especially in convex landscapes, but it requires repeated free energy calculations via MBAR at each epoch, resulting in longer runtimes, and its uncertainty estimates are less reliable in non-convex landscapes. Overall, posterior sampling is faster and more robust, while variational minimization may be preferable for very high-dimensional, convex problems.
@@
This representation ensures continuity across the periodic boundary at $\pm\pi$ and provides a smooth input feature space for learning.
@@
The network was trained using ADAM optimization for 2000 epochs with a learning rate of $10^{-3}$, using the Good-Bad likelihood model with the BICePs score as the loss function. The toy dataset consisted of 5 conformational states, each with 60 synthetic $J$-coupling observables using the Karplus relation $J(\phi)=A\cos^{2}(\phi)+B\cos(\phi)+C$, with ground truth coefficients $A=6.51$, B=1.76B=-1.76, and $C=1.60$. Our training dataset consisted of 300 $\phi$-angles as NN inputs.
First, we test the performance without adding random or systematic noise to the synthetic experimental J-coupling observables. Figure S16a shows the convergence behavior of the BICePs score across five independent training runs, where a plateau is reached around step 1500. Figure S16b shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). For comparison, we also include the result of singular value decomposition (SVD) applied directly to the Karplus form under the assumption that backbone $\phi$ angles are known precisely, as would be the case with a high-resolution crystal structure. The parameters from SVD were determined to be $A=6.47\pm 0.002$, $B=-1.75\pm 0.001$, and $C=1.62\pm 0.001$. Our framework yields comparable predictive performance, where the predictive accuracy of J-couplings for the SVD approach achieved an RMSE of $0.01\pm 0.001$ Hz, while the BICePs-trained neural network reached an RMSE of $0.04\pm 0.01$ Hz.
Next, we evaluate the performance when adding random or systematic noise to the synthetic experimental $J$-coupling observables ($\sigma_{\text{data}}=0.79$ Hz). Figure S17a–b shows the convergence behavior of the BICePs score and gradient norm across five independent training runs, where convergence begins around 1000 epochs. The BICePs score was determined to be $f_{\xi=0\rightarrow 1}=62.29$ nats. Figure S17c shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). Comparison against SVD illustrates that BICePs, when used to train neural networks from conformational ensembles alone, recovers predictive performance comparable to analytical fitting based on ground-truth dihedral angles, even in the presence of random and systematic errors. When comparing predictive accuracy of $J$-couplings, the SVD approach achieved an RMSE of $0.24\pm 0.03$ Hz, while the BICePs-trained neural network reached an RMSE of $0.21\pm 0.01$ Hz, demonstrating that our framework yields improved predictive performance. For completeness, we also used BICePs to optimize the parameters of the Karplus relation in this example, obtaining $A=6.6\pm 0.05$, $B=-1.7\pm 0.023$, and $C=1.7\pm 0.037$, which resulted in the lowest RMSE of $0.17\pm 0.0001$ Hz. This outcome is expected, since the “true” values were originally generated using the Karplus relation. By contrast, the neural network achieves remarkably low error without being constrained to any functional form, highlighting its flexibility and strong predictive capability.
@@
Despite these advantages, neural networks trained on limited or biased datasets can still face overfitting risks. Several mitigation strategies can be incorporated within the BICePs training framework: (1) data augmentation via additional synthetic observables or sampling of underrepresented dihedral regions, (2) regularization through dropout layers to discourage over-parameterization, and (3) physically informed constraints on network architecture or inputs, such as torsional symmetries—as we did for J-coupling or local environment descriptors for chemical shifts. These strategies, combined with the adaptive regularization provided by marginalizing over $\sigma$, further safeguard against overfitting and improve model generalizability.
@@
Models have limited predictive reliability in regions lacking sufficient training data. A central advantage of our framework lies in its ability to incorporate physically relevant information through the choice of input features or explicit regularization. In our example using the toy model system, we employed the backbone dihedral angle $\phi$ as the sole input feature (projected to the unit circle to maintain periodicity). However, neural network inputs can encompass a broader range of features for complex observables such as chemical shifts, including quantum calculations, partial charges, descriptors of local atomic environments, interatomic distances, hydrogen bonding patterns, solvent accessibility, or torsional strain energies. These diverse features encode crucial structural and electronic contexts and can be precomputed from molecular simulations or quantum calculations before training.
@@
We did not impose physical constraints or priors when training our model to predict $J$-coupling constants. Users might explicitly incorporate such physical priors or constraints through architectural designs, regularization terms, or physically-informed input features, especially critical when dealing with sparse or unevenly distributed training data. For instance, in our experiments, neural networks employing Rectified Linear Unit (ReLU) activations struggled to accurately predict $J$-couplings at backbone dihedral angles greater than $90^{\circ}$. In contrast, GELU activations provided smoother and more reliable extrapolation in these underrepresented regions. Thus, choosing suitable activation functions and rich input features improves predictive accuracy across the conformational landscape.
Relation to mixture-based outlier models and multi-level outlier severity.
Our outlier-aware Good-and-Bad likelihood model uses a two-component mixture $M=2$ of variances. One down-weights data affected by systematic error, while the other preserves the influence of "good" data. A natural extension of the Good-Bad model replaces a single "bad" mode with multiple levels of outlier severity.
In reliability and survival analysis, censored finite-mixture models help separate heterogeneous data while accounting for censoring. For example, Akhtar & AlharthiAkhtar and Alharthi (2025) apply a censored two-component mixture of geometric distributions under a Bayesian framework, illustrating how finite mixtures capture central trends while disregarding extreme values. Although distinct from our context, it illustrates the broader principle that finite mixtures model structured heterogeneity rather than ad hoc data exclusion.
For a natural extension to our Good-Bad model, we assign $p(\sigma_{j}|\sigma_{0})$ from equation 7 to be a multi-severity Good–Bad prior
$p(\sigma_{j}\mid\sigma_{0},\{\omega_{m}\},\{\varphi_{m}\})=\sum_{m=1}^{M}\omega_{m}\,\delta\big(\sigma_{j}-\varphi_{m}\sigma_{0}\big),$ (23)
where $\omega_{m}\geq 0$ are mixture weights with $\sum_{m=1}^{M}\omega_{m}=1$, and the severity inflators satisfy $\varphi_{1}\equiv 1<\varphi_{2}<\cdots<\varphi_{M}$. Thus, $\varphi_{m}$ scales the typical uncertainty $\sigma_{0}$ to represent increasingly severe outliers (mild $\to$ severe). The standard two-component Good–Bad model is recovered at $M=2$.
Inserting equation 23 into the integrand of equation 7 gives, for each observable $j$,
$\int\limits_{\sigma^{\mathrm{SEM}}}^{\infty}$ $\mathcal{N}\big(d_{j}\mid g_{j},\sigma_{j}^{2}\big)\,\sum_{m=1}^{M}\omega_{m}\,\delta(\sigma_{j}-\varphi_{m}\sigma_{0})\,d\sigma_{j}$ (24)
$=\sum_{m=1}^{M}\omega_{m}\,H\big(\varphi_{m}\sigma_{0}-\sigma^{\mathrm{SEM}}\big)\,\mathcal{N}\big(d_{j}\mid g_{j},\varphi_{m}^{2}\sigma_{0}^{2}\big),$
i.e., each observation is evaluated under a finite Gaussian scale mixture with the Heaviside factor enforcing the $\sigma^{\mathrm{SEM}}$ lower bound.
If we integrate out the mixing weights $\omega_{m}$} as in our two-component case, a uniform Dirichlet prior that is flat on the simplex $\omega_{m}\geq 0,\ \sum_{m}\omega_{m}=1$} implies $\mathbb{E}[\omega_{m}]=1/M$ by symmetry, leading to
$\sum_{m=1}^{M}$ $\left[\int\limits_{\Delta^{M-1}}\omega_{m}\,p(\bm{\omega})\,d\bm{\omega}\right]\,H\big(\varphi_{m}\sigma_{0}-\sigma^{\mathrm{SEM}}\big)\mathcal{N}\big(d_{j}\mid g_{j},\varphi_{m}^{2}\sigma_{0}^{2}\big)$ (25)
$=\frac{1}{M}\sum_{m=1}^{M}H\big(\varphi_{m}\sigma_{0}-\sigma^{\mathrm{SEM}}\big)\,\mathcal{N}\big(d_{j}\mid g_{j},\ \varphi_{m}^{2}\sigma_{0}^{2}\big).$
Practically, this multi-severity formulation leaves a single Bayesian uncertainty parameter $\sigma_{0}^{B}$ and introduces M1M-1 free severity parameters $\varphi_{m}$} (with $\varphi_{1}=1$ fixed) to be inferred in the posterior. This extends the single-severity Good–Bad model used in our framework (where only a single $\varphi$ is sampled alongside $\sigma_{0}^{B}$) and provides finer control over how aggressively outliers are down-weighted.
@@ V Conclusion
The BICePs algorithm is openly available at github.com/vvoelz/biceps. All calculations in this work were performed using the development version biceps_v3.0a, which is planned to be merged into the main branch. The specific version can be accessed at: github.com/vvoelz/biceps/tree/biceps_v3.0a. All scripts, Jupyter notebook examples and analysis can be found here. Detailed documentation and tutorials can be found here: biceps.readthedocs.io. For any issues or questions, please submit the request on GitHub.
@@ Acknowledgements
The authors acknowledge that substantial content in this manuscript has also been included in the doctoral dissertation of RMR.Raddi (2024) RMR, TM and VAV are supported by National Institutes of Health grant R01GM123296. This research includes calculations carried out on HPC resources supported in part by the National Science Foundation through major research instrumentation grant number 1625061 and by the US Army Research Laboratory under contract number W911NF-16-2-0189. The authors also thank Kim Sharp for valuable discussions on extending the Good-and-Bad likelihood framework to incorporate multiple modes of data severity, and Somaiyeh Dadashi for her assistance with illustrations.
@@
[…]
@@
Appendix B: Method equivalence between joint posterior sampling and variational minimization.
Equivalence of objectives.
We define the BICePs score in the main text (Eq. 15), but rewrite it here for clarity. Let $Z(\theta)$ be the model evidence given a set of parameters against a well-defined constant reference $Z_{0}$, then
$f(\theta)=-\ln\left(Z(\theta)\big/Z_{0}\right),$ (S5)
where
$Z(\theta)=\iint\exp\left(-u(\mathbf{X},\mathbf{\sigma};\theta)\right)d\mathbf{X}d\mathbf{\sigma},$ (S6)
and the unnormalized log-density
$u(X,\sigma;\theta)\ =\ -\ln\!\big(p(D|X,\sigma,\theta)p(X)p(\sigma)\big)+C,$ (S7)
for some constant $C$.
The joint posterior (Eq. 3) can be written as
$p(X,\sigma,\theta|D)\propto\ p(\theta)\,\exp\!\big(-u(X,\sigma;\theta)\big),$ (S8)
Marginalizing $(X,\sigma)$ yields
$p(\theta\mid D)$ $\propto\ p(\theta)\,Z(\theta)$ (S9)
$\implies-\ln p(\theta\mid D)$ $=\ f(\theta)\ -\ \ln p(\theta)\ +C.$
Therefore, maximizing $p(\theta|D)$ during joint posterior sampling is equivalent to minimizing the BICePs Score $f(\theta)$ (when applying a flat prior for $p(\theta)$).
Equivalence of gradients.
Rewritting Eq. 17 as
$\begin{split}\frac{\partial f(\theta)}{\partial\theta_{i}}&=\iint\frac{1}{Z(\theta)}\left[\frac{\partial u}{\partial\theta_{i}}\right]\exp\left(-u\right)d\mathbf{X}d\mathbf{\sigma}=\bigg\langle\frac{\partial u}{\partial\theta_{i}}\bigg\rangle\end{split}$ (S10)
The joint posterior sampling approach can use $M$ uncorrelated MCMC samples $(X^{(m)},\sigma^{(m)})$ to approximate the first derivative of the BICePs score:
$\frac{\partial f(\theta)}{\partial\theta_{i}}\ \approx\frac{1}{M}\sum_{m=1}^{M}\frac{\partial u(X^{(m)},\sigma^{(m)})}{\partial\theta_{i}}$ (S11)
@@ Methods
Table S1: Coefficients for the Karplus relation $^{3}\!J(\phi)=A\cos^{2}(\phi+\phi_{0})+B\cos(\phi+\phi_{0})+C$, determined by SVD on the 1D3Z ensemble using uniform populations.
$\phi_{0}$ $\mathrm{A}$ (Hz) $\mathrm{B}$ (Hz) $\mathrm{C}$ (Hz)
${{}^{3}\!J}_{C^{\prime}C}$ $0^{\circ}$ $0.55\pm 0.29$ $0.94\pm 0.04$ $0.740\pm 0.04$
${{}^{3}\!J}_{C^{\prime}C^{\beta}}$ $60^{\circ}$ $2.32\pm 0.73$ $0.12\pm 1.07$ $0.13\pm 0.36$
${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$ $120^{\circ}$ $4.23\pm 0.07$ $2.24\pm 0.09$ $1.03\pm 0.05$
${{}^{3}\!J}_{H^{N}C^{\prime}}$ $180^{\circ}$ $4.87\pm 0.40$ $1.36\pm 0.10$ $0.05\pm 0.05$
${{}^{3}\!J}_{H^{N}C^{\beta}}$ $60^{\circ}$ $3.04\pm 0.12$ $0.21\pm 0.14$ $0.35\pm 0.05$
${{}^{3}\!J}_{H^{N}H^{\alpha}}$ $60^{\circ}$ $7.93\pm 0.16$ $1.78\pm 0.07$ $0.85\pm 0.14$
Table S2: Mean absolute error (MAE) for SVD(1D3Z) and BICePs(1D3Z) parameter sets across three different ensembles. These MAE values for BICePs(1D3Z) are found in Figures S13-S15.
1D3Z 2NR2 CHARMM22*
SVD(1D3Z) BICePs(1D3Z) SVD(1D3Z) BICePs(1D3Z) SVD(1D3Z) BICePs(1D3Z)
${{}^{3}\!J}_{C^{\prime}C}$ 0.11 0.11 0.17 0.16 0.10 0.09
${{}^{3}\!J}_{C^{\prime}C^{\beta}}$ 0.15 0.14 0.22 0.22 0.20 0.19
${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$ 0.23 0.23 0.27 0.27 0.29 0.29
${{}^{3}\!J}_{H^{N}C^{\prime}}$ 0.29 0.29 0.35 0.35 0.34 0.33
${{}^{3}\!J}_{H^{N}C^{\beta}}$ 0.19 0.19 0.24 0.24 0.25 0.25
${{}^{3}\!J}_{H^{N}H^{\alpha}}$ 0.41 0.41 0.62 0.61 0.72 0.70
@@
Figure S2: BICePs predicted forward model parameters in the presence of random and systematic error ($\sigma_{data}=0.471$ Hz) for a toy model system. (a) Karplus curves predicted for SVD (orange) and BICePs (blue), where the "true" (black dashed line) parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$. The extracted parameters from the SVD fitting were found to be $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$ and the BICePs was $\{A=6.6\pm 0.037,B=-1.8\pm 0.016,C=1.5\pm 0.027\}$, averaged over three independent chains. The uncertainty is represented by the thickness of the curves. When comparing the predicted J-couplings to the True J-couplings, the SVD approach yield an RMSE of $0.15\pm 0.02$ Hz and BICePs gives an RMSE of $0.06\pm 0.007$ Hz. For the BICePs calculation, we used the Good-Bad likelihood model with 32 replicas and burned for 20k steps, followed by 50k steps of additional sampling. Red dots correspond to the syntetic experimental J-coupling data points. (b) The marginal posterior distribution of uncertainty $p(\sigma_{J})$. The maximum a posteriori was determined to be $\hat{\sigma}_{J}=0.272$ Hz, and the a posteriori variance scaling parameter $\hat{\varphi}_{J}=1.98$. (c) Landscapes of the BICePs score, $f$ for pairs of Karplus coefficients. (d) The marginal posterior distribution of Karplus coefficients.
Figure S3: The Student’s model gives similar performance to the Good-Bad model. Karplus coefficients predicted using the Student’s likeliood model ($\{A=6.8\pm 0.033,B=-1.9\pm 0.025,C=1.4\pm 0.033\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$ Hz). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2. When comparing the predicted J-couplings to the True J-couplings, the SVD approach yield an RMSE of $0.15\pm 0.02$ Hz and BICePs gives an RMSE of $0.14\pm 0.011$ Hz.
@@
Figure S10: Traces of sampled Karplus coefficients for the 1d3z ensemble over 50k steps of $MC$, post-burn. BICePs calculations used the Good-Bad model with 32 replicas. Traces display low variance with no jumps, which demonstrates converged samples
@@
Figure S16: BICePs trains NN parameters for $J$-coupling predictions using a toy model system. (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1500 training steps. (b) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters A=6.51,B=1.76,C=1.60{A=6.51,B=-1.76,C=1.60}. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters $A=6.47\pm 0.002,B=-1.75\pm 0.001,C=1.62\pm 0.001$}. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. No random or systematic noise was added to the synthetic data. Red dots correspond to the syntetic experimental J-coupling data points. When comparing the predicted J-couplings to the True J-couplings, the SVD approach yield an RMSE of $0.01\pm 0.001$ Hz and BICePs trained NN predictions have an RMSE of $0.04\pm 0.01$ Hz.
Figure S17: BICePs trains NN parameters for $J$-coupling predictions using a toy model system in the presence of random and systematic error ($\sigma_{\text{data}}=0.79$ Hz). (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1000 training steps. (b) The average gradient norm across the five chains during training. (c) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters A=6.51,B=1.76,C=1.60{A=6.51,B=-1.76,C=1.60}. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters $A=6.39\pm 0.13,B=-1.55\pm 0.06,C=1.94\pm 0.10$}. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. Red dots correspond to the syntetic experimental $J$-coupling data points. When comparing the predicted J-couplings to the True J-couplings, the SVD approach yield an RMSE of $0.24\pm 0.03$ Hz and BICePs trained NN predictions have an RMSE of $0.21\pm 0.01$ Hz.